Pesto Company possesses 80 percent of Salerno Company's outstanding voting stock. Pesto uses the initial value method to account for this investment. On January 1, 2014, Pesto sold 7 percent bonds payable with a $11.5 million face value (maturing in 20 years) on the open market at a premium of $650,000. On January 1, 2017, Salerno acquired 40 percent of these same bonds from an outside party at 96.6 percent of face value. Both companies use the straight-line method of amortization. For a 2018 consolidation, what adjustment should be made to Pesto's beginning Retained Earnings as a result of this bond acquisition?

  • $361,700 increase

  • $368,200 increase

  • $377,400 increase

  • $355,200 increase
Answer #1

Book Value = 11,500,000 + 650,000 = 12,150,000

Amortization = 650,000/20 x 3 = 97,500

Cash Received = 11,500,000 x 40% x 96.6%

=4,443,600

Book Value = 12,150,000 - 97,500 = 12,052,500

Book Value of Retired Bonds = 4,821,000

Cash Received = 4,443,600

Gain on Retirement of bonds = 377,400

Cash Interest Expense = 11,500,000 x 40% x 7% = 322,000

Premium Amortization = 650,000/20 x 40% = 13,000

Interest Expense = 322,000 - 13,000 = 309,000

Discount amortization = 11,500,000 x 40%/20 - 3 Years x (100% - 96.6%)

=9,200

=322,000 + 9,200 = 331,200

Adjustment = 377,400 + 309,000 - 331,200

=355,200

So Increase is D. 355,200 Increase